Chao, who is an experienced architect, can draw a certain set of plans in 6 hours. It takes Carl, who is a new architect, 10 hours to draw the same set of plans. How long would it take them working together to draw the set of plans?
Answer by checkley77(12844)

6*10/(10+6)=60/16=3.75 HOURS WORKING TOGETHER.